determine cos theta to three decimal places where sin theta=6/square root 61 and theta is an angle in the second quadrant


1
Expert's answer
2021-10-07T14:37:50-0400

"Sin\space θ
=\frac{6}{\sqrt{61}}\\Sin^2\space θ
+
Cos^2\space θ
=1\\Cos^2\space θ
=1-Sin^2\space θ
\\Cos\space θ
=\sqrt{1-Sin^2\space θ
}\\\\Cos\space θ
=\sqrt{1-Sin^2 \space θ
}\\
Cos\space θ
=\sqrt{1-(\frac{6}{\sqrt{61}})^2}\\
Cos\space θ
=\sqrt{1-\frac{36}{61}}\\

Cos\space θ
=\sqrt{\frac{61-36}{61}}=\frac{\sqrt{25}}{\sqrt{61}}\\

Cos\space θ
=\pm \frac{5}{\sqrt{61}}"

second quadrant "Cos\space θ" is negative

"=-\frac{5}{7.810}\\


Cos\space θ
=-0.640"